OpenStudy (anonymous):

Simplify 1/4 (g − 16)

OpenStudy (anonymous):

g/4 - 4

OpenStudy (anonymous):

1/4g - 4

OpenStudy (anonymous):

g/4- 4

myininaya (myininaya):

Nullify are you understanding what they are doing? I'm just curious because you post alot of similar questions?

OpenStudy (anonymous):

No I don't /:

OpenStudy (anonymous):

you have to distribute the 1/4 to the g and 16.

OpenStudy (anonymous):

\[a(b+c) = ab +ac\]

myininaya (myininaya):

\[a(b+c)=ab+ac \text{ <- this is called the distribution property}\]